A, B, C and D went fishing. A caught 50% more fish than B. If A gives C 20 fish, A, B and C
will have the same number of fish. A, B and C together caught 4 times more fish than D. They
kept 20% of the fish they caught, sold the rest at 200 sum per fish and shared the money equally.
How much money did each take home?

Answer by Theo(13342)

A, B, C, and D went fishing.
let a = amount of fish A caught.
let b = amount of fish B caught.
let c = amount of fish C caught.
let d = amount of fish D caught.

A caught 50% more fish than B.
This means that a = b + .5b which means that a = 1.5b.

If A gives 20 fish to C, then A, B, and C will have the same number of fish each.
This means that a - 20 = b = c + 20.

A + B + C together caught 4 times more fish than D.
This means that a + b + c = 4d.

we know that a = 1.5b
we also know that a - 20 = b
if we substitute 1.5b for a in the second equation, we get:
1.5b - 20 = b
subtract b from both sides of this equation and add 20 to both sides of this equation to get:
1.5b - b = 20
simplify to get:
.5b = 20
divide both sides of this equation by b to get:
b = 40

since a = 1.5b, this means that:
a = 60

since c + 20 = b, this mean that:
c + 20 = 40
subtract 20 from both sides of this equation to get:
c = 20

we now have:
a = 60
b = 40
c = 20

we also know that a + b + c = 4d.
a + b + c = 120
120 = 4d
divide both sides of this equation by 4 to get:
120/4 = d which means that:
d = 30

we now have:
a = 60
b = 40
c = 20
d = 30

the total is 150.
they keep 20% which means they keep 30.
that means they sold 120 at 200 apiece.
that means they received 120 * 200 = 24000
they divided it equally which means they divided 24000 by 4 to get 6000 apiece.

that last statement could possibly be interpreted differently.
the statement i am talking about is:
"A, B and C together caught 4 times more fish than D."

my own wording for that statement would have been:
"A, B, and C together caught 4 times as many fish as D."
That is the interpretation I used.
My answer is based on that interpretation.

The second interpretation would lead to a different answer.

the statement:
"A, B, and C, together, caught 4 times more fish than D." can be interpreted as follows:

They caught 4 times more fish than D.
This means they caught 5 times as many fish as D caught.
If D catches 1 fish and they catch 5 fish, then they caught 4 times more fish than D because 5 minus 1 is equal to 4 which is 4 times 1.
If D catches 2 fish and they catch 10 fish, then they caught 4 times more fish than D because 10 minus 2 = 8 which is 4 times 2.

Under this alternate interpretation, the answer would be changed as follows:
starting from "we also know that":
we also know that a + b + c = 4d + d which is equal to 5d.
a + b + c = 120
120 = 5d
divide both sides of this equation by 5 to get:
120/5 = d which means that:
d = 24
this means that:
d = 24
a + b + c = 5d = 120
120 - 24 = 96
96 = 4 * 24.
they caught 5 times more fish than D caught.

we now have:
a = 60
b = 40
c = 20
d = 24

the total is 144.
they keep 20% which means they keep 28.8
that means they sold 115.2 at 200 apiece.
that means they received 115.2 * 200 = 23040
they divided it equally which means they divided 23040 by 4 to get 5760 apiece.

this alternate interpretation does not divide as evenly as the first interpretation.
that makes it a little more doubtful that this interpretation is what they meant.

I believe the wording of the statement:
"A, B and C together caught 4 times more fish than D"
should have been worded as, and may have been meant to be worded as:
"A, B, and C together caught 4 times as many fish as D caught."

I recommend to go with the original solution because I think that's what they meant.